# Predicting Maintenance Cessation of Open Source Software Repositories with An Integrated Feature Framework

### v1

@@ 4. Feature Framework
Numerous factors contribute to the deaths of open source projects (Robbes et al., 2012; Sawant et al., 2016). We design a feature framework capturing multiple perspectives in OSS maintenance, including surface activity, user participation, maintainer behavior, and project evolution patterns. Our features are constructed from large-scale, longitudinal repository data, organized into three main categories: user-centric, maintainer-centric, and project evolution features. In the following subsections, we describe each group and its computation.
The remainder of this section systematically describes the design and computation of each feature category, and empirically justifies their inclusion via ablation and interpretability analysis in Section 4.2.
4.1. Feature Design
We first summarize the surface features commonly adopted in existing literature and then motivate our enhancements via new structured feature groups. Table 2 provides an overview of all major feature categories used in our framework.
Table 2. Categorized list of features used in the cessation prediction framework.
@@ 4.1. Feature Design
contrib_diversity Contributor diversity measured by 1 - Gini index.
balance_index Structural balance combining maintainer and others contribution share.
@@ 4.1.1. Surface Features
Prior work on OSS health and abandonment prediction has largely relied on surface-level features. These include cumulative or interval-based counts of stars, commits, issues, PRs, tags, and comments—all of which are directly accessible via public platform APIs and have intuitive appeal as signals of project activity or visibility (Liao et al., 2023; Moid et al., 2021; Robinson et al., 2022; Chen et al., 2014).
While these features provide a useful starting point, they exhibit several critical limitations in our setting:
Lack of user differentiation: Surfacer features treat all user actions equally, overlooking differences in user roles. For example, a star from a renowned developer is more meaningful than one from a novice, but surface features flatten such distinctions.
Vulnerability to manipulation: Surface features are easily inflated (e.g., via paid stars or bots (He et al., 2024)), undermining their reliability for predicting project health.
Limited expressive power: Surface features blur the line between maintainers and casual users, and cannot capture collaboration quality or project lifecycle changes—leading to missed signals about maintenance trajectories.
To address these shortcomings and move beyond the limitations of surface features, we next introduce three structured feature categories.
@@ 4.1.2. User-Centric Features
To overcome the limitations of surface activity features, which do not distinguish user roles and are prone to manipulation, we introduce user-centric features based on the structure and evolution of user-repository interactions. Our key insight is that users attracted to high-quality projects tend to be more experienced, and repositories that engage these users are more likely to be sustainable. Some studies have uncovered new information within networks (Clauset et al., 2005; Pinzger et al., 2008; Nagappan et al., 2010). We represent the OSS ecosystem as a weighted bipartite graph (Blincoe et al., 2016) (Figure 3) and iteratively propagate influence between users and repositories. This approach enables us to capture complex patterns of user influence and project health that surface features alone cannot reveal.
@@
In our formulation, each user-repository pair is linked through various types of interactions, including starring, committing, forking, and issues. Building upon the intuition that the quality and timing of user engagement are more indicative of sustainability than sheer quantity, we simply assign interaction-specific base weights: $w_{orig}^{star}=1$, $w_{orig}^{commit}=8$, $w_{orig}^{fork}=4$, $w_{orig}^{issue}=2$. To further emphasize early or persistent engagement, a temporal decay factor is applied to each interaction, defined as:
(1) $p_{(up)}^{t}=\frac{k_{(up)}^{t}}{N_{p}},\quad D(p)=\frac{1}{1+p},\quad d_{(up)}^{t}=D(p_{(up)}^{t})$
[…]
@@
To better quantify the technical contribution of commits, we additionally weight each commit by the logarithm of lines of code changed ($LOC$):
@@
This user-centric formulation offers several advantages. By taking into account the historical quality of user participation, it distinguishes habitual contributors from one-off actors and thus uncovers subtle signals of project health. Additionally, the graph-based design reduces susceptibility to manipulation, as artificially inflated surface features do not result in higher network-based influence. Finally, the incorporation of temporal weighting highlights sustained or early engagement, which is closely associated with long-term project sustainability.
@@
Similar to many key features in software engineering (Goeminne and Mens, 2011; Zhang et al., 2021), our proposed user-centric features also exhibit a pronounced long-tail distribution. To ensure feature comparability across different time periods and repositories, we apply two normalization strategies: (1) percentile ranking, which transforms scores into relative ranks within each time window (Mujahid et al., 2022), and (2) Z-score normalization, with logarithmic transformation to address skewness:
@@ 4.1.3. Maintainer-Centric Features
Maintainers hold primary responsibility for project evolution, governance, and the formal declaration of maintenance cessation, their behavioral signals serve as salient predictors of project risk. To this end, we design maintainer-centric features that directly track maintainer engagement and responsiveness within the repository lifecycle.
@@
An important feature of governance continuity is the time elapsed since the last maintainer activity. For each repository and timestamp, we compute the maintainer inactivity interval as the duration since the most recent repository-level action performed by any identified maintainer. This inclusive approach captures the latest evidence of maintainer presence, beyond privileged operations like merges or direct commits. A prolonged inactivity interval strongly signals disengagement and is frequently a precursor to maintenance cessation. Compared to user activity features, this feature provides finer granularity in distinguishing actual maintenance cessation from superficial reductions in community activity.
@@
Healthy open source projects typically exhibit prompt maintainer responses to externally submitted PRs and issues, indicating ongoing stewardship and community integration. As the risk of cessation grows, response times elongate or become absent altogether. We capture this dynamic through two related features:
@@
Contribution Ratio and Community Balance.
Stability in the proportion of maintainer involvement relative to overall project actions is a further signal of sustainable governance. Over-reliance on maintainers (or, conversely, abrupt decline in their participation) may indicate community imbalance or capacity exhaustion. We therefore include as features:
  • Maintainer Contribution Ratio: The fraction of all substantial actions (commits, merges, issues) performed by maintainers in a sliding window.
  • Imbalance Feature: Quantified as the squared deviation from an empirical ideal ratio (e.g., 0.5), emphasizing both under- and over-concentration of maintainer activity.
@@ 4.1.4. Project Evolution Features
Beyond user and maintainer dynamics, the long-term evolution of an OSS project reveals critical signals of sustainability, community health, and risk of maintenance cessation. To capture these deeper aspects, we construct a suite of project evolution features that model not only activity levels but also the changing collaboration structure, development focus, and periodic behavioral patterns over the lifecycle of a repository.
@@
Mature open source projects typically follow an “onion model” comprising core maintainers, active contributors, and peripheral users. Disruption to the balance among these roles—such as disproportionate maintainer dominance or a decline in contributor diversity—often precedes project maintenance cessation. We operationalize this via:
@@
  • Participation Imbalance ($d$): The squared deviation from an ideal maintainer ratio: $d=(p-0.5)^{2}$. Lower values indicate a healthy, balanced distribution of responsibilities, while high values signal imbalance and increased cessation risk.
@@
The relative focus of project development—whether on adding new features or on bug fixing—naturally shifts as a repository matures. A transition from innovation-driven contributions to predominantly bug-fixing is often symptomatic of declining momentum: PRs are classified as feature-oriented or bug-fix based on curated keyword matching in PR titles (e.g., “feature”, “add” for features; “fix”, “bug” for bug-fixes). For each time window, we extract: the ratio of feature to bug-fix PRs, as well as the individual proportions of each category. A shrinking feature/bug-fix PR ratio is an early marker of functional stagnation and elevated risk of upcoming maintenance cessation.
@@
A consistent pattern of negative deviation serves as a quantifiable warning signal preceding cessation events.
@@ 4.2. Feature Effectiveness Validation
Just as physiological indicators can predict health outcomes, behavioral traces of OSS repositories can serve as “biomarkers” for project vitality. In our setting, the interval from repository creation to maintenance cessation constitutes the software “lifespan.” Survival analysis is well suited for this task, as it effectively handles time-to-event data with extensive right-censoring—matching the characteristics of our dataset. In this section, we evaluate the utility of our proposed features.
@@
Survival analysis is particularly apt for this problem because: (1) OSS projects may not all undergo cessation within the observation window, resulting in right-censored data. (2) Survival models naturally support event-time analysis, enabling either risk ranking (“which projects are likely to cessation sooner?”) or direct estimation of maintenance cessation time.
Rather than predicting a point estimate of cessation time, survival models rank project risk according to concordance with observed lifespans. The concordance index (C-index) is therefore the primary evaluation metric, measuring the fraction of predicted risk orderings that match actual event orderings. It ranges from 0.5 (random guessing) to 1.0 (perfect discrimination). Empirically, a C-index above 0.8 is considered strong for prediction (Liao et al., 2023).
@@
We instantiate the AFT approach using XGBoost’s survival module (Cox, 2018; Avinash Barnwal and Hocking, 2022), where the log of event time (maintenance lifespan) is regressed on input features.
The AFT model supports efficient optimization with right-censored samples using negative log-likelihood loss (nloglik). We partition the data into 80% training and 20% test splits, with early stopping and hyperparameter. We observe rapid convergence and stable loss minimization of the AFT (XGBoost-based) model in training.
@@
We observe that augmenting the surface features (B) with any individual feature group consistently improves predictive performance. Notably, user-centric features (U) yield the largest gain, increasing Harrel’s C-index from 0.748 to 0.826—surpassing the effect of adding HITS scores (H) alone (0.810). Furthermore, using only U, M, and P—excluding all surface features—achieves near-parity with the full model (0.838 vs. 0.846), demonstrating the sufficiency and complementarity of the proposed features.
@@
[…]
@@
The results show that User-Centric Features contribute most prominently to the model, highlighting the predictive value of user engagement patterns (Ait et al., 2022). Maintainer-Centric Features follow with a substantial score of 2213.24, suggesting the importance of maintainer activity signals. Project Evolution Features also exhibit meaningful predictive strength, capturing long-term development dynamics. In contrast, Surface Features contribute the least, indicating limited standalone utility in cessation risk modeling.
@@
The beeswarm SHAP visualization demonstrates that latest_ maintainer_activity_interval, weight_zscore, and contrib_diversity emerge as the three most influential predictors. These 3 features representing each of our novel feature categories, consistently exhibit the strongest directional impact on model predictions, confirming their critical role in cessation predicting.
@@
To illustrate the advantages of our newly engineered features, we examine the time-series evolution of multiple signals from the Electronic WeChat project (Zhongyi Tong et al., [n. d.]) (Figure 5). All feature values are min-max normalized to the [0.1, 1] range for a beautiful view.
@@
This project, a popular third-party MacOS WeChat client, ceased maintenance after Tencent officially released WeChat for Mac 2.0 on August 16, 2016 and updates regularly thereafter.
@@
Surface Feature (Orange): The number of stars remained consistently high before the cessation event, reflecting the project’s lasting popularity. However, this persistent star growth results mainly from historical reputation and user recommendations, making it insensitive to changes in actual maintenance or activity. This indicates star count does not provide a timely warning of approaching maintenance cessation.
User-centric Feature (Blue): The percentile rank of interaction weight exhibits a steady rise as cessation approaches. This increase indicates a decline in genuine community engagement, even while stars remain high. As fewer influential users interact meaningfully with the project, this feature effectively signals a loss of active user support, often preceding maintenance termination.
Maintainer-centric Feature (Green): The time since the last maintainer activity increases continuously before maintenance cessation, showing that maintainers are gradually reducing their participation. This growing inactivity interval quantitatively foreshadows the decline of maintainer commitment, aligning with the actual cessation event.
Project Evolution Feature (Red): Initially, the maintainer contribution ratio is high but then drops sharply before cessation. This shift reflects a breakdown in the stable collaboration pattern, signaling impending inactivity.
This case underscores the critical need for advanced, multi-perspective features—beyond traditional surface features—when predicting OSS maintenance cessation with precision and interpretability. Our three newly proposed feature types capture a broader range of repository state information than surface activity features.

### v2

@@ 4. Feature Framework
Numerous factors contribute to the abandonment of OSS projects (Robbes et al., 2012; Sawant et al., 2016). To model this phenomenon comprehensively, we design a multi-perspective feature framework. Our framework integrates several feature categories, beginning with surface activity features which establish a baseline using common features (e.g., stars, commits) widely used in prior work. Moving beyond this foundation, we introduce three novel feature groups that constitute our primary contribution: (1) user-centric features—to our best knowledge, the first of their kind proposed for abandonment prediction, modeling the OSS ecosystem as a weighted bipartite graph to capture the structure and evolution of user-repository interactions; (2) maintainer-centric features, whose novelty lies in explicitly identifying maintainers—a critical distinction from generic contributors often glossed over in studies—leveraging their specific engagement and responsiveness signals for prediction; and (3) project evolution features, which mark another first by operationalizing into a predictive model deeper lifecycle characteristics such as the changing collaboration structure, development focus (e.g., the bug-fix ratio (Soto et al., 2016; Tufano et al., 2019)), and periodic behavioral patterns. In the following subsections, we delineate the design and computation for each feature group and explain how each captures behavioral mechanisms relevant to program comprehension, with empirical validation provided in Section 4.2. We next describe each group and their relevance to program comprehension in Section 4.1, with empirical tests in Section 4.2.
Table 2. Categorized list of features used in the abandonment prediction framework.
@@
contrib_diversity Contributor diversity measured by the 1-Gini index.
balance_index Structural balance combining maintainer and others’ contribution share.
@@
4.1. Feature Design
We begin by summarizing surface features commonly used in existing work, then introduce our enhanced (new) feature groups. Table 2 provides an overview of all features in our framework.
@@ 4.1.1. Surface Features
Prior work mainly uses surface features: counts of stars, commits, issues, PRs, etc., all of which are directly accessible via public platform APIs and have intuitive appeal as signals of project activity (Liao et al., 2023; Moid et al., 2021; Robinson et al., 2022; Chen et al., 2014; Coelho and Valente, 2017; Borges et al., 2016).
While intuitive, these features suffer from key limitations: (1) No user differentiation: all interactions are weighted equally, overlooking differences in user roles. For example, a star from a renowned developer is more meaningful than one from a novice, but surface features flatten such distinctions; (2) Vulnerability to manipulation: they can be artificially inflated (e.g., via paid stars (He et al., 2024)), undermining their reliability for predicting project health; (3) Limited expressiveness: they fail to capture community roles, collaboration patterns, and project evolution.
To address these shortcomings from a program comprehension perspective, we propose three structured feature categories that move beyond surface counts to capture interpretable behavioral and structural signals.
@@ 4.1.2. User-Centric Features
We introduce user-centric features based on the structure and evolution of user-repository interactions—a novel perspective not previously explored in abandonment prediction. Our key insight is that users attracted to high-quality projects tend to be more experienced, and repositories that engage these users are more likely to be sustainable. Some studies have uncovered new information within networks (Clauset et al., 2005; Pinzger et al., 2008; Nagappan et al., 2010). We represent the OSS ecosystem as a weighted bipartite graph (Blincoe et al., 2016) (Figure 3), where nodes represent users or repositories and timestamped edges capture their interactions. Figure 3 serves as a schematic of this structure, and details are provided in this subsection. This approach enables us to capture complex patterns of user influence and project health that surface features alone cannot reveal.
@@
We model user-repository relationships through four interaction types: starring, committing, forking, and issues. Reflecting that engagement quality and timing better indicate sustainability than volume (supported by the superior C-index of S+U over S+H in Table 3), we assign base weights to each interaction type. Given the dataset’s size, which precludes exhaustive tuning, we determined these weights ($w_{orig}^{star}=1$, $w_{orig}^{commit}=8$, $w_{orig}^{fork}=4$, $w_{orig}^{issue}=2$) by evaluating multiple combinations on a 500-repository subset. This configuration performed well on the full dataset, and further parameter optimization is secondary to our primary contribution: establishing the overall effectiveness of user-centric features for abandonment prediction and understanding the status of OSS projects. To prioritize early or sustained engagement, we apply a temporal decay factor to each interaction, defined as:
(1) $pos_{(up)}^{t}=\frac{k_{(up)}^{t}}{N_{p}^{t^{\prime}}},\quad D(pos)=\frac{1}{1+p},\quad d_{(up)}^{t}=D(pos_{(up)}^{t})$
Here, $pos_{(up)}^{t}$ denotes the relative position of the interaction between user $u$ and OSS project (repository) $p$ at time $t$ with second-level granularity; $k_{(up)}^{t}$ indicates that the action is the $k$-th occurrence between $u$ and $p$ in their timeline; and $N_{p}^{t^{\prime}}$ is the total number of actions associated with $p$ at the target time for calculating the data. $D(pos)$ is a generic decay function that can be instantiated as needed, and $d_{(up)}^{t}$ represents the temporal decay factor for the action at $(u,p,t)$, assigning higher weights to earlier interactions to reflect the user’s foresight.
@@
To better quantify the technical contribution of commits, we additionally weight each commit by the logarithm of lines of code changed $LOC$, a metric derived from commit payloads. Due to API limitations, $LOC$ excludes non-file changes, but we argue this has minimal impact on the user-repository network, as active users typically also make file modifications.
@@
This user-centric formulation offers several advantages. By taking into account the historical quality of user participation, it distinguishes habitual contributors from one-off actors and thus uncovers subtle signals of project health. Additionally, the graph-based design reduces susceptibility to manipulation, as artificially inflated surface features do not result in higher network-based influence. Further, the incorporation of temporal weighting highlights sustained or early engagement, which is closely associated with long-term project sustainability. Ultimately, this yields a more sturdy insight into project attention and maintenance status than what surface features can reveal.
From a program comprehension standpoint, these user influence signals directly support dependency vetting and ecosystem analysis by providing context on the quality of the user base for each project. Understanding which projects attract and retain influential users aids developers in evaluating the reliability and trustworthiness of their dependencies during comprehension, onboarding, or refactoring activities.
@@
Our user-centric features exhibit a heavy-tailed distribution, typical in software engineering (Goeminne and Mens, 2011; Zhang et al., 2021). To enable cross-project and temporal comparison, we apply two normalization strategies: (1) percentile ranking within each time window (Mujahid et al., 2022), and (2) Z-score normalization on log-transformed values to mitigate skewness:
@@ 4.1.3. Maintainer-Centric Features
Maintainers hold primary responsibility for project evolution, governance, and the formal declaration of abandonment; consequently, their behavioral signals serve as salient predictors of project risk. However, prior studies primarily rely on repository-level activity intervals (e.g., consecutive inactive days) derived from commit logs, which overlook this critical human factor (Coelho et al., 2020; Decan et al., 2020). Furthermore, some studies attempted role categorization without real community roles and did not achieve high C-index (Valiev et al., 2018). To address these gaps, we propose a novel set of Maintainer-Centric Features that directly track maintainer engagement and responsiveness. Our approach moves beyond these limitations by first explicitly identifying the maintainers and then modeling their behavioral signals, providing a more direct and powerful predictor of abandonment risk.
In addition to their role in abandonment prediction, maintainer-centric features directly assist program comprehension tasks. For example, knowledge of maintainer activity intervals and response patterns enables developers to gauge project stewardship, identify potential bottlenecks in collaboration, and understand governance structures—key factors when making maintenance, onboarding, or technical debt management decisions.
@@
An important feature of governance continuity is the time elapsed since the last maintainer activity. For each repository and timestamp, we compute the maintainer inactivity interval as the duration since the most recent repository-level action performed by any identified maintainer. This inclusive approach captures the latest evidence of maintainer presence, beyond privileged operations like merges or direct commits. A prolonged inactivity interval strongly signals disengagement and is frequently a precursor to abandonment.
@@
Healthy open source projects typically exhibit prompt maintainer responses to externally submitted PRs and issues, indicating ongoing stewardship and community integration. As the risk of abandonment grows, response times elongate or become absent altogether. We capture this dynamic through two related features:
@@ 4.1.4. Project Evolution Features
Beyond user and maintainer dynamics, the long-term evolution of an OSS project reveals critical signals of sustainability, community health, and risk of abandonment. To capture these deeper aspects, we construct a suite of project evolution features that build upon and extend prior concepts. For instance, while prior studies (Soto et al., 2016; Tufano et al., 2019) have discussed the bug-fix ratio, we are the first to systematically incorporate it, along with the changing collaboration structure, development focus, and periodic behavioral patterns, into a predictive model for abandonment.
For program comprehension, these evolution features help developers and architects grasp the shifting focus, maturity, and collaborative structure of a project over time. By interpreting transitions in feature versus bug-fix focus, contributor diversity, and deviations from historical behavior, stakeholders can more accurately assess project trajectories, anticipate maintenance challenges, and plan sustainable architecture or integration strategies.
@@
Mature open source projects typically follow an “onion model” comprising core maintainers, active contributors, and peripheral users (Nakakoji et al., 2002). Disruption to the balance among these roles—such as disproportionate maintainer dominance or a decline in contributor diversity—often precedes project abandonment. We operationalize this via:
@@
  • Participation Imbalance ($d$): The squared deviation from an ideal maintainer ratio: $d=(p-0.5)^{2}$. Lower values indicate a healthy, balanced distribution of responsibilities, while high values signal imbalance and increased abandonment risk.
@@
The relative focus of project development—whether on adding new features or on bug fixing—naturally shifts as a repository matures. A transition from innovation-driven contributions to predominantly bug-fixing is often indicates declining momentum: PRs are classified as feature-oriented or bug-fix based on curated keyword matching in PR titles (e.g., “feature”, “add” for features; “fix”, “bug” for bug-fixes). For each time window, we extract: the ratio of feature to bug-fix PRs, as well as the individual proportions of each category. The shrinking feature/bug-fix PR ratio is an early marker of functional stagnation and elevated risk of upcoming abandonment.
@@
A consistent pattern of negative deviation serves as a quantifiable warning signal preceding abandonment events.
@@ 4.2. Feature Effectiveness Validation
Just as physiological indicators can predict health outcomes, behavioral traces of OSS repositories can serve as “biomarkers” for project vitality. In our setting, the interval from repository creation to abandonment constitutes the software “lifespan.” Survival analysis is well suited for this task, as it effectively handles time-to-event data with extensive right-censoring—matching the characteristics of our dataset. In this section, we evaluate the utility of our proposed features.
@@
Survival analysis is particularly apt for this problem because: (1) OSS projects may not all undergo abandonment within the observation window, resulting in right-censored data. (2) Survival models naturally support event-time analysis, enabling either risk ranking (“which projects are likely to be abandoned sooner?”) or direct estimation of abandonment time.
Rather than predicting a point estimate of abandonment time, survival models rank project risk according to concordance with observed lifespans. The concordance index (C-index) is therefore the primary evaluation metric, measuring the fraction of predicted risk orderings that match actual event orderings; it ranges from 0.5 (random guessing) to 1.0 (perfect discrimination). Empirically, a C-index above 0.8 is considered strong for prediction (Liao et al., 2023).
@@
We instantiate the AFT approach using XGBoost’s survival module (Avinash Barnwal and Hocking, 2022; Zhong et al., 2021; Alvarez et al., 2025; Cox, 2018), where the log of event time (maintenance lifespan) is regressed on input features.
The AFT model supports efficient optimization with right censored samples using negative log-likelihood loss (nloglik). When applying the AFT model, it is standard practice to randomly split the dataset into training and testing sets (Avinash Barnwal and Hocking, 2022; Zhong et al., 2021; Alvarez et al., 2025). In line with this convention, we randomly partitioned our data into 80% for training and 20% for testing, and employed early stopping during the model training to prevent overfitting. We observe rapid convergence and stable loss minimization of the AFT (XGBoost-based) model in training.
@@
We observe that augmenting the surface features (S) with any individual feature group consistently improves predictive performance. Notably, user-centric features (U) yield the largest gain, increasing Harrel’s C-index from 0.748 to 0.826—surpassing the effect of adding HITS scores (H) alone (0.810). Furthermore, using only U, M, and P—excluding all surface features—achieves near-parity with the full model (0.838 vs. 0.846), demonstrating the sufficiency and complementarity of the proposed features.
@@
To elucidate the relative contribution of different feature groups, we leverage XGBoost’s built-in F-score metric, which reflects how frequently a feature is used to split nodes across the ensemble (Chen and Guestrin, 2016). For each feature group, we aggregate the F-scores of its constituent features using a geometric mean (i.e., $\exp(\mathrm{mean}(\log(\text{F-score})))$), which smooths the influence of extreme values and provides a more balanced importance estimate across heterogeneous features. Figure 4 visualizes the aggregated scores for each group.
@@
The results show that User-Centric Features contribute most prominently to the model, highlighting the predictive value of user engagement patterns (Ait et al., 2022). Maintainer-Centric Features follow with a substantial score of 2213.24, suggesting the importance of maintainer activity signals. Project Evolution Features also exhibit meaningful predictive strength, capturing long-term development dynamics. In contrast, Surface Features contribute the least, indicating limited standalone utility in abandonment risk modeling.
@@
The beeswarm SHAP visualization demonstrates that latest_ maintainer_activity_interval, weight_zscore, and contrib_diversity emerge as the three most influential predictors. These 3 features representing each of our novel feature categories, consistently exhibit the strongest directional impact on model predictions, confirming their critical role in abandonment predicting.
@@
To illustrate the advantages and interpretability of our newly engineered features, we examine the time-series evolution of multiple signals from the Electronic WeChat project (Zhongyi Tong et al., [n. d.]). All feature values are min-max normalized to the [0.1, 1] range for a beautiful view.
@@
This project, a popular third-party MacOS WeChat client, was abandoned after Tencent officially released WeChat for Mac 2.0 on August 16, 2016, which has been updating regularly ever since.
@@
Surface Feature (Orange): The number of stars remained consistently high before the abandonment event, reflecting the project’s lasting popularity. However, this persistent star growth results mainly from historical reputation and user recommendations, making it insensitive to changes in actual maintenance or activity. This observation aligns with our earlier discussion in Section 4.1.1, confirming that the inherent limitations of surface metrics like star count, limit their effectiveness as a timely warning signal for project abandonment.
User-centric Feature (Blue): The percentile rank of interaction weight exhibits a steady increase (indicating a decline in relative ranking) as abandonment approaches. This change indicates a decline in genuine community engagement, even while stars remain high. As fewer influential users interact meaningfully with the project, this feature effectively signals a loss of capable user support, often preceding maintenance termination.
Maintainer-centric Feature (Green): The time since the last maintainer activity increases continuously prior to abandonment, confirming a complete cessation of involvement. This extended inactivity interval serves as a quantitative indicator that the project has been neglected by its maintainers, directly foreshadowing the eventual abandonment event.
Project Evolution Feature (Red): Initially, the maintainer contribution ratio is high but then drops sharply before abandonment. This shift reflects a breakdown in the stable collaboration pattern, signaling impending inactivity.
This case underscores the critical need for advanced, multi-perspective features—beyond traditional surface features—when predicting OSS abandonment with precision and interpretability. Our three newly proposed feature types capture a broader range of repository state information than surface activity features. Such case studies demonstrate how our multi-perspective feature set can offer not only predictive alerts but also actionable narratives for project stakeholders to interpret and comprehend the abandonment process in real time.
Comprehensive Feature Efficacy
Collectively, the ablation studies, F-Score analysis, SHAP, and case analysis provide converging evidence that all feature categories summarized in Table 2 contribute effectively to abandonment prediction. Together, these multi-perspective features improve prediction accuracy and give developers interpretable signals for understanding project health, collaboration, and evolution, supporting both risk assessment and program comprehension.